# The differences of baryonic and dark matter scaling relations in galaxy clusters between IllustrisTNG simulations and observations

### v1

@@ 2.2 Galaxy clusters in observations
To compare our results, we used the observations of C18, based on 91 galaxy clusters identified through the Sunyaev and Zeldovich effect (SZ) (Sunyaev & Zeldovich, 1972) within the redshift range $0.25\leq z\leq 1.25$. This sample was selected from the SPT-SZ survey (Bleem et al., 2015) and matched with Chandra X-ray observations. These measurements were further correlated with optical photometry in the griz bands from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016) and near-infrared photometry obtained using either the Wide-field Infrared Survey Explorer (WISE; Wright et al., 2010) or the Infrared Array Camera (IRAC; Fazio et al., 2004) aboard the Spitzer telescope.
C18 used the $\zeta$-$M_{500}$ scaling relations for the SZ-effect observable provided by the SPT collaboration, the $\zeta$ normal distribution with the SZ-effect signal-to-noise ratio $\xi$ as the unit width, and the calibrations from de Haan et al. (2016) to estimate the total mass $M_{\rm 500c}$ for each cluster in the sample (based on Bocquet et al., 2015). The mass range of these clusters is $M_{\rm 500c}=\left(2.63-12.71\right)\times 10^{14}$ $\rm M_{\odot}$. The hot-gas mass, $M_{\rm hot-gas,500c}$, was determined using Chandra X-ray observations by fitting the best-fitting modified $\beta$-model (Vikhlinin et al., 2009) out to $R_{\rm 500c}$ (based on McDonald et al., 2013). The stellar mass, $M_{\rm star,500c}$, was estimated by fitting the spectral energy distributions (SEDs) of individual galaxies, constructing a stellar mass function (SMF; Schechter, 1976), and summing the stellar masses of all galaxies in the cluster. Finally, the baryon fractions were calculated as the sum of the X-ray gas mass and the stellar mass of galaxies, normalized by $M_{\rm 500c}$. The baryon fractions for these clusters range from $f_{\rm Chiu,500c}\approx 0.07-0.21$.
@@ 3 Scaling relations
Most of this section is based on the mathematical developments of Kaiser (1986), Mulroy et al. (2019), and P22. We follow their mathematical framework for the self-similar scaling relations of baryonic fractions in Subsection 3.1. Finally, the fitting procedures for the scaling relations are defined in Subsection 3.2.
@@ 3.1 Self-similar model and baryonic mass fraction scaling relations
We assume that the hot gas is in hydrostatic equilibrium. Thus, we expect the pressure, $P_{\rm gas}$, and the density, $\rho_{\rm gas}$, of the gas to satisfy Eq. (1).
$\dfrac{1}{\rho_{\rm gas}}\dfrac{{\rm d}P_{\rm gas}}{{\rm d}r}=-\dfrac{GM}{r^{2}}$ (1)
The solution to the previous equation for the total mass, as a function of the gas density $\rho_{\rm gas}(r)$ and temperature $T(r)$ profiles within the cluster, is given in Eq. (2).
$M(<r)=-\dfrac{k_{B}rT(r)}{\mu m_{p}G}\left(\dfrac{{\rm d\ln}\rho_{\rm gas}(r)}{{\rm d\ln}r}+\dfrac{{\rm d\ln}T(r)}{{\rm d\ln}r}\right)$ (2)
If the mass is defined based on a spherical collapse model, the mass enclosed within a spherical region of radius $R_{\Delta}$, with $\Delta$ times the critical density of the Universe $\rho_{\rm c}(z)$, is given by Eq. (3).
$M_{\Delta}=\dfrac{4\pi}{3}\Delta\rho_{\rm c}(z)R_{\Delta}^{3}$ (3)
@@
$\rho_{\rm c}(z)=\dfrac{3H(z)^{2}}{8\pi G}$ (4)
It is important to mention that the definitions of $R_{\Delta}$ and $M_{\Delta}$ depend on the overdensity value $\Delta$. In this study, we consider $\Delta=\{500,200\}$.
In the self-similar model (Kaiser, 1986), the properties of baryonic matter are scaled solely by the total cluster mass, independent of redshift. Thus, $M_{\rm star,\Delta}\propto M_{\rm gas,\Delta}\propto M_{\rm star,\Delta}+M_{\rm gas,\Delta}=M_{\rm bar,\Delta}\propto M_{\Delta}$. We define the fraction of a given baryonic component $\rm X$ (stellar, gas, or baryonic = stellar + gas) as shown in Eq. (5).
$f_{\rm X,\Delta}\equiv M_{\rm X,\Delta}/M_{\Delta}$ (5)
The scaling relation for these fractions is given in Eq. (6).
$f_{\rm star,\Delta}\propto f_{\rm gas,\Delta}\propto f_{\rm bar,\Delta}\propto M_{\Delta}^{0}=1$ (6)
@@ 3.2.1 Simple Power Law
The self-similarity model assumes that the scaling relations of galaxy clusters follow a simple power law (SPL: Lotka, 1926) distribution (Kaiser, 1986). If we have two observables, $(X,Y)$, the simple power law model relates them using two free parameters, $(A_{\rm SPL},\alpha_{\rm SPL})$. Eq. (7) presents the simple power law model.
$Y=A_{\rm SPL}X^{\alpha_{\rm SPL}}$ (7)
@@ 3.2.2 Broken Power Law and General Double Power Law
One of the findings from P22 was that a simple power law is insufficient for describing the scaling relations in IllustrisTNG simulations. Feedback processes are more intense in low-mass haloes and groups of galaxies (Ayromlou et al., 2021), breaking the self-similarity of the scaling relations for these low-mass haloes. To address this, we use the General Double Power Law (GDPL) model. This model is based on the broken power law (BPL: Jóhannesson et al., 2006), where the data is divided into two regions, separated by a pivot point, $X_{\rm pivot}$. This value determines the slope values: $\alpha_{\rm BPL,1}$ for $X\leq X_{\rm pivot}$ and $\alpha_{\rm BPL,2}$ for $X>X_{\rm pivot}$. In the GDPL model, the transition between the slopes $\alpha_{\rm GDPL,1}$ and $\alpha_{\rm GDPL,2}$ is smoothed by a parameter $\delta$, that quantifies the width of the transition region between the slopes (Pop et al., 2022a).
The broken power law (BPL) model is defined as,
@@
The inclusion of a factor $X_{\rm pivot}^{-1}$ in the pieces ensures that they coincide at $X=X_{\rm pivot}$, thereby guaranteeing continuity at that point.
The general double power law (GDPL) model is defined as,
$Y=A_{\rm GDPL}\left(\frac{X}{X_{\rm pivot}}\right)^{\alpha_{\rm GDPL,1}}\left\{\frac{1}{2}\left[1+\left(\frac{X}{X_{\rm pivot}}\right)^{\frac{1}{\delta}}\right]\right\}^{\left(\alpha_{\rm GDPL,2}-\alpha_{\rm GDPL,1}\right)\delta}$ (9)
The main advantage of the GDPL model over the SPL model is that it allows us to identify where self-similarity in the scaling relations holds. The $X_{\rm pivot}$ value defines the breakpoint between the first and second slopes, delineating the mass range where self-similarity breaks. While adding multiple slopes with several breakpoints could improve the model fitting performance, a single $X_{\rm pivot}$ with two slopes is sufficient to capture the effects of AGN feedback within the mass range of our sample, as shown in P22.
@@ 4.1 Comparison between observations and IllustrisTNG
C18 measured halo masses using the SZ-effect, hot gas using the X-ray $\beta$-model, and the stellar mass of each galaxy through SED fitting and SMF construction, all within $R_{\rm 500c}$. Thus, we consider two measurements within the simulations:
@@
  • We took only the gas from the cluster with $T\geq 10^{7}$ K enclosed within a sphere of radius $R_{\rm 500c}$. Regarding the stellar mass, it corresponds to the sum of the stellar masses of galaxies within a spherical aperture of twice the stellar half-mass radius ($2\times R_{M_{\rm star}/2}\sim r_{\rm op}$). Warm gas, cold gas, and the intracluster light (ICL) are not considered in this analysis.
  • All gas and stars within the halo within a sphere of radius $R_{\rm 500c}$. This includes hot gas, warm gas, cold gas, galaxy stars, and the intracluster light (ICL).
@@
Figure 1: Distribution of the percentage difference between the stellar masses from subfind and $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ apertures at different constrains: in an halocentric radius $R_{\rm 500c}$ (black), $R_{\rm 200c}$ (red), and the mass considered by the friends-of-friends (FoF) algorithm (blue). The dashed lines represent the normal distribution fitted in their respective constraint (and their respective color).
@@
The first approach is essential for comparing TNG100 and TNG300 with the observations from C18. The second helps examine the behaviour of the scaling relations when all components expected to be in the clusters are considered. We do not use the subfind aperture, as our goal is to create a measurement procedure as similar as possible to the observations and to characterize all baryonic components within the clusters. Fig. 1 shows the differences between the stellar masses defined by a subfind and a $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ apertures within $R_{\rm 500c}$, $R_{\rm 200c}$, and FoF constraints regarding stellar masses. We find that the differences are approximately $(31.88\pm 1.94)$ per cent for TNG100 and $(32.53\pm 2.74)$ per cent for TNG300 within $R_{\rm 500c}$, comparable to those observed for other constraints. These differences are significant when calculating baryon fractions, fitting the SPL scaling relation, and comparing them with the scaling relations from C18. In the case of masses within $R_{\rm 200c}$ and FoF, the percentages remain unchanged relative to $R_{\rm 500c}$, showing a tendency to range between $30.92\%$ and $32.21\%$ in both simulations.
@@
Figure 2: The Simple Power Law (SPL) fit (green region) of the IllustrisTNG haloes (black dots), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$, in a radius $R_{\rm 500c}$, and the redshift $z$. All panels have their corresponding simulation name label and a value for the $\alpha_{\rm SPL}$ slope from the SPL fit in the upper-left corner. Top panels: The fraction of baryons includes the hot-gas and the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. Bottom panels: The fraction of baryons includes all the gas and the subfind aperture. The shade areas are represented with a width of $3\sigma$.
@@
Furthermore, we test the self-similarity of IllustrisTNG haloes. This aims to remove the redshift dependence from the baryon fraction, allowing us to fit a SPL that depends solely on the mass, as described by Eq. (6). We calculate the baryon fraction using Eq. (5) for haloes from TNG100 and TNG300 inside an $R_{500}$ as described above. The SPL used in this case is the relation $f_{\rm bar,500c}\propto\left(1+z\right)^{\alpha_{\rm SPL}}$. Fig. 2 shows the SPL fits of the IllustrisTNG haloes for the relation between baryon fraction and redshift, considering both comparison types. In the first type of comparison, the slopes of the SPL are $\alpha_{\rm SPL}\simeq 0.029\pm 0.051$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.071\pm 0.021$ for TNG300. In the second type of comparison, the slopes are $\alpha_{\rm SPL}\simeq 0.006\pm 0.023$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.037\pm 0.010$ for TNG300. These slopes were obtained using a non-linear least squares regression fit to our SPL model. The covariance matrix of the fit determines the slope errors. In the self-similar model presented by Kaiser (1986), the theoretical slope should be close to zero, as the baryonic matter should only scale with the total cluster mass, independently of redshift. Examining the slope values closely, all are positive, indicating a slight change in the scaling relations, with a minimal increase in the baryon fraction with redshift. However, the differences between the self-similar and simulation slopes are small, on the order of hundredths (and, in one case, thousandths). Thus, we confirm that the simulation slopes are effectively zero, removing the redshift dependence from the baryon fraction. Thus, the baryon fraction $f_{\rm bar,500c}$ is expected to scale only with the mass $M_{\rm 500c}$.
@@
Figure 3: The Simple Power Law (SPL) fit (green region) of the IllustrisTNG haloes (black dots), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$ and halo mass $M$ in a radius $R_{\rm 500c}$, for all redshifts. All panels have their respective simulation tag and a value for the $\alpha_{\rm SPL}$ slope from the SPL fit in the upper-left corner. Top panels: The fraction of baryons includes the hot-gas and the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. Bottom panels: The fraction of baryons includes all the gas and all the stars inside the halo. The yellow areas represent the Planck Collaboration et al. (2016a) fraction of baryons. Also, the C18 data (blue dots) is shown with its SPL fit (blue areas). The shade areas are represented with a width of $3\sigma$.
@@
Using the same baryon fractions, we perform an SPL fit for IllustrisTNG haloes to determine the slopes, $\alpha_{\rm SPL}$, and compare them with the data from C18. The SPL used in this case is the relation $f_{\rm bar,500c}\propto\left(M_{\rm 500c}\right)^{\alpha_{\rm SPL}}$. Fig. 3 presents these results alongside the Universe’s baryon fraction measured by Planck Collaboration et al. (2016a). The top panels clearly show that the observable data does not directly match TNG100 or TNG300. The reason for this discrepancy is that IllustrisTNG has an AGN feedback model that enriches efficiently the intra-cluster medium gas of the low massive clusters and the galaxy groups (Pop et al., 2022a). As a consequence, the fraction of baryons is raised at low mass values. However, this does not impact our analysis, as we focus on the behavior of the scaling relation. In this context, we find that in IllustrisTNG, the relations between the baryon fraction and halo mass yield a slope $\alpha_{\rm SPL}\simeq 0.248\pm 0.018$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.261\pm 0.006$ for TNG300. Without considering redshift, C18 data provides an SPL scaling relation with a slope of $\alpha_{\rm Chiu}\simeq 0.350\pm 0.076$. The scaling relations for TNG100, TNG300, and C18 data show similar behavior, with baryon fractions increasing at high halo masses and decreasing at low halo masses. Furthermore, the slope values from IllustrisTNG and C18 agree within $1-2\sigma$. This agreement suggests that the IllustrisTNG results may provide insights into the missing baryons problem and help reconcile the tension with the hierarchical scenario highlighted by C18.
@@
In the bottom panels of Fig (3), we show how the behavior of the scaling relations changes when the remaining gas ($T<10^{7}$ K) and all remaining stellar mass in the halo are included. At first glance, a direct comparison with C18 is not possible, as the parameters are not directly comparable. Although the general behavior remains, showing an increase in the baryon fraction, the slopes differ significantly between simulations and observations. We find a slope $\alpha_{\rm SPL}\simeq 0.071\pm 0.010$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.120\pm 0.003$ for TNG300 in the scaling relations. These lower slopes arise from including the remaining gas and accounting for the ICL in the measurements, which benefits low-mass haloes more than higher-mass ones. Consequently, low-mass haloes exhibit a greater fraction of cold gas compared to high-mass haloes (see Sections 4.2 and A). Additionally, more haloes in both simulations reach the universal fraction of baryon. Most haloes that fail to reach this value have masses of $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}\lesssim 14$. This mass threshold aligns with values reported in numerous studies discussing the breakdown of self-similarity in scaling relations (Pop et al., 2022a) and the redistribution of baryons in low-mass haloes by non-gravitational processes, such as AGN feedback (Ayromlou et al., 2023).

### v2

@@ 2.2 Galaxy clusters in observations
To compare our results, we used the observations of C18, which analysed 91 galaxy clusters identified via the SZ effect(Sunyaev & Zeldovich, 1972) in the redshift range $0.25\leq z\leq 1.25$. This sample was selected from the SPT-SZ survey (Bleem et al., 2015) and matched with Chandra X-ray observations. These measurements were further correlated with optical photometry in the griz bands from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al., 2016) and near-infrared photometry obtained using either the Wide-field Infrared Survey Explorer (WISE; Wright et al., 2010) or the Infrared Array Camera (IRAC; Fazio et al., 2004) on the Spitzer telescope.
C18 used the $\zeta$-$M_{500}$ scaling relations for the SZ effect observable, provided by the SPT collaboration, and the $\zeta$ normal distribution with the signal-to-noise ratio of the SZ effect $\xi$ as the unit width. The calibrations of de Haan et al. (2016) were used to estimate the total mass, $M_{\rm 500c}$, for each cluster in the sample (based on Bocquet et al., 2015). The mass range of these clusters was $M_{\rm 500c}=\left(2.63-12.71\right)\times 10^{14}$ $\rm M_{\odot}$. The hot-gas mass, $M_{\rm hot-gas,500c}$, was determined using Chandra X-ray observations by fitting the best-fitting modified $\beta$-model (Vikhlinin et al., 2009) out to $R_{\rm 500c}$ (based on McDonald et al., 2013). The stellar mass, $M_{\rm star,500c}$, was estimated by fitting the SEDs of individual galaxies, constructing a stellar mass function (SMF; Schechter, 1976), and summing the stellar masses of all galaxies in the cluster. Finally, the baryon fractions were calculated as the sum of the X-ray gas mass and the stellar mass of galaxies, normalised by $M_{\rm 500c}$. The baryon fractions for these clusters ranged from $f_{\rm Chiu,500c}\approx 0.07-0.21$.
@@ 3 Scaling relations
This section mainly uses the mathematical frameworks of Kaiser (1986), Mulroy et al. (2019), and P22. We followed their mathematical framework for the self-similar scaling relations of baryonic fractions in Subsection 3.1. Finally, the fitting procedures for the scaling relations are defined in Subsection 3.2.
@@ 3.1 Self-similar model and baryonic mass fraction scaling relations
We assumed that the hot gas was in hydrostatic equilibrium. Thus, we expected the pressure, $P_{\rm gas}$, and the density, $\rho_{\rm gas}$, of the gas to satisfy Eq. (1):
$\dfrac{1}{\rho_{\rm gas}}\dfrac{{\rm d}P_{\rm gas}}{{\rm d}r}=-\dfrac{GM}{r^{2}},$ (1)
where the solution to the previous equation for the total mass, as a function of the gas density $\rho_{\rm gas}(r)$ and temperature $T(r)$ profiles within the cluster, is given in Eq. (2):
$M(<r)=-\dfrac{k_{B}rT(r)}{\mu m_{p}G}\left(\dfrac{{\rm d\ln}\rho_{\rm gas}(r)}{{\rm d\ln}r}+\dfrac{{\rm d\ln}T(r)}{{\rm d\ln}r}\right).$ (2)
If the mass is defined on the basis of a spherical collapse model, the mass enclosed within a spherical region of radius $R_{\Delta}$, with $\Delta$ times the critical density of the Universe $\rho_{\rm c}(z)$ can be derived. This is given by Eq. (3):
$M_{\Delta}=\dfrac{4\pi}{3}\Delta\rho_{\rm c}(z)R_{\Delta}^{3}.$ (3)
@@
$\rho_{\rm c}(z)=\dfrac{3H(z)^{2}}{8\pi G}.$ (4)
It is important to mention that the definitions of $R_{\Delta}$ and $M_{\Delta}$ depend on the overdensity value $\Delta$. In this study, we considered $\Delta=\{500,200\}$.
In the self-similar model (Kaiser, 1986), the properties of baryonic matter are scaled solely by the total cluster mass, independent of redshift. Thus, $M_{\rm star,\Delta}\propto M_{\rm gas,\Delta}\propto M_{\rm star,\Delta}+M_{\rm gas,\Delta}=M_{\rm bar,\Delta}\propto M_{\Delta}$. We defined the fraction of a given baryonic component $\rm X$ (stellar, gas, or baryonic = stellar + gas) as shown in Eq. (5):
$f_{\rm X,\Delta}\equiv M_{\rm X,\Delta}/M_{\Delta}.$ (5)
The scaling relation for these fractions is given in Eq. (6):
$f_{\rm star,\Delta}\propto f_{\rm gas,\Delta}\propto f_{\rm bar,\Delta}\propto M_{\Delta}^{0}=1,$ (6)
@@ 3.2.1 Simple power law
The self-similarity model assumes that the scaling relations of galaxy clusters follow a simple power-law (SPL) distribution (SPL: Lotka, 1926) (Kaiser, 1986). If we have two observables, $(X,Y)$, the SPL model relates them using two free parameters, $(A_{\rm SPL},\alpha_{\rm SPL})$. Eq. (7) presents the simple power-law model:
$Y=A_{\rm SPL}X^{\alpha_{\rm SPL}}.$ (7)
@@ 3.2.2 Broken power law and general double power law
One of the findings of P22 was that an SPL is insufficient to describe the scaling relations in IllustrisTNG simulations. Feedback processes are more intense in low-mass haloes and groups of galaxies (Ayromlou et al., 2021), breaking the self-similarity of the scaling relations for these low-mass haloes. To address this, we used the general double power-law (GDPL) model. This model is based on the broken power law (BPL: Jóhannesson et al., 2006), in which the data are divided into two regions, separated by a pivot point $X_{\rm pivot}$. This value determines the slope values: $\alpha_{\rm BPL,1}$ for $X\leq X_{\rm pivot}$ and $\alpha_{\rm BPL,2}$ for $X>X_{\rm pivot}$. In the GDPL model, the transition between the slopes $\alpha_{\rm GDPL,1}$ and $\alpha_{\rm GDPL,2}$ is smoothed by a parameter $\delta$, which quantifies the width of the transition region between the slopes (Pop et al., 2022a).
The BPL model was defined as
@@
where the inclusion of a factor $X_{\rm pivot}^{-1}$ in both pieces ensures that the two branches coincide at $X=X_{\rm pivot}$, thereby guaranteeing continuity at the pivot point. The GDPL model was defined as
$Y=A_{\rm GDPL}\left(\frac{X}{X_{\rm pivot}}\right)^{\alpha_{\rm GDPL,1}}\left\{\frac{1}{2}\left[1+\left(\frac{X}{X_{\rm pivot}}\right)^{\frac{1}{\delta}}\right]\right\}^{\left(\alpha_{\rm GDPL,2}-\alpha_{\rm GDPL,1}\right)\delta},$ (9)
with the main advantage of the GDPL model over the SPL model being that it allowed us to identify where self-similarity in the scaling relations holds. The parameter $X_{\rm pivot}$ defined the breakpoint between the first and second slopes, delineating the mass range where self-similarity breaks. Although introducing additional slopes with multiple breakpoints could improve the model’s fitting performance, we found that a single $X_{\rm pivot}$ with two slopes was sufficient to capture the effects of AGN feedback within the mass range of our sample, as demonstrated in P22.
@@ 4.1 Comparison between observations and IllustrisTNG
C18 measured halo masses using the SZ-effect, hot gas using the X-ray $\beta$-model, and the stellar masses of each galaxy through SED fitting and SMF construction, all within $R_{\rm 500c}$. Thus, to enable a direct comparison, we considered two measurement approaches within the simulations. First, we included only the gas from the cluster with $T\geq 10^{7}$ K enclosed within a sphere of radius $R_{\rm 500c}$. The stellar mass in this case was the sum of the stellar masses of galaxies within a spherical aperture of twice the stellar half-mass radius ($2\times R_{M_{\rm star}/2}\sim r_{\rm op}$). In this approach, warm gas, cold gas, and the ICL were not considered. Second, we considered all gas and stars within the halo inside a sphere of radius $R_{\rm 500c}$. This included hot gas, warm gas, cold gas, galaxy stars, and the ICL.
@@
Figure 1: Distribution of the percentage difference between the stellar masses from Subfind and $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ apertures under different constrains. The black, red, and blue lines correspond to measurements within halocentric radius $R_{\rm 500c}$, $R_{\rm 200c}$, and the mass considered by the FoF algorithm, respectively. The dashed lines represent the normal distribution fitted to each constraint and are shown in their corresponding colour.
@@
The first approach was essential for comparing TNG100 and TNG300 with the observations of C18. The second approach enabled examination of the scaling relation behaviour when all components expected to be in the clusters are included. We did not use the Subfind aperture, as our goal was to create a measurement procedure that closely matches the observations and to characterise all baryonic components within the clusters. Fig. 1 shows the differences between the stellar masses defined by Subfind and $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ apertures within $R_{\rm 500c}$, $R_{\rm 200c}$, and FoF constraints. For TNG100, the difference was approximately $31.88\pm 1.94$ per cent, and for TNG300, $32.53\pm 2.74$ per cent within $R_{\rm 500c}$, comparable to values observed for other constraints. These differences are significant for calculating baryon fractions, fitting the SPL scaling relation, and comparing with the scaling relations of C18. For masses within $R_{\rm 200c}$ and FoF, the percentages remained unchanged relative to $R_{\rm 500c}$, with values ranging from $30.92\%$ and $32.21\%$ in both simulations.
@@
Figure 2: SPL fit of baryon mass fraction $f_{\rm bar}$ in IllustrisTNG at $R_{\rm 500c}$ and redshift $z$. Black dots show individual haloes; the green region indicates the SPL fit. Each panel is labelled with their corresponding simulation tag and the value for the $\alpha_{\rm SPL}$ slope from the SPL fit in the upper-left corner. Top panels: The fraction of baryons includes the hot gas and the galaxy stellar mass within the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. Bottom panels: The fraction of baryons includes all the gas and all stars inside the halo. The shaded areas are represented with a width of $3\sigma$.
@@
Next, we tested the self-similarity of IllustrisTNG haloes. This was intended to remove the redshift dependence from the baryon fraction, allowing us to fit an SPL that depended solely on the mass, as described by Eq. (6). We calculated the baryon fraction using Eq. (5) for TNG100 and TNG300 haloes within a $R_{500}$ as described above. The SPL used in this case was the relation $f_{\rm bar,500c}\propto\left(1+z\right)^{\alpha_{\rm SPL}}$. Fig. 2 shows the SPL fits of the IllustrisTNG haloes for the relation between baryon fraction and redshift, considering both comparison types. In the first comparison, the slopes were $\alpha_{\rm SPL}\simeq 0.029\pm 0.051$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.071\pm 0.021$ for TNG300. In the second type of comparison, the slopes were $\alpha_{\rm SPL}\simeq 0.006\pm 0.023$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.037\pm 0.010$ for TNG300. These slopes were obtained using a non-linear least squares regression fit to our SPL model. The errors on the slopes were determined from the covariance matrix of the fit. In the self-similar model presented by Kaiser (1986), the theoretical slope should be close to zero, as the baryonic matter should only scale with the total cluster mass, independently of redshift. Upon examining the slope values closely, we found that all were positive, indicating a slight change in the scaling relations, with a minimal increase in the baryon fraction with redshift. However, the differences between the self-similar and simulation slopes were small, on the order of hundredths (and in one case, thousandths). We therefore confirmed that the simulation slopes were effectively zero, removing the redshift dependence from the baryon fraction. As a result, the baryon fraction $f_{\rm bar,500c}$ was expected to scale only with the mass $M_{\rm 500c}$.
@@
Figure 3: SPL fit of baryon mass fraction $f_{\rm bar}$ in IllustrisTNG at $R_{\rm 500c}$ for all redshifts. Black dots show individual haloes; the green region indicates the SPL fit. Each panel is labelled with their corresponding simulation tag and the value for the $\alpha_{\rm SPL}$ slope from the SPL fit in the upper-left corner. Top panels: The fraction of baryons includes the hot gas and the galaxy stellar mass within the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. Bottom panels: The fraction of baryons includes all the gas and all the stars inside the halo. The yellow areas represent the Planck Collaboration et al. (2016a) fraction of baryons. The C18 data (blue dots) is also shown, with its SPL fit (blue areas). The shaded areas are represented with a width of $3\sigma$.
@@
Using the same baryon fractions, we performed an SPL fit for IllustrisTNG haloes to determine the slopes, $\alpha_{\rm SPL}$, and compared them with the data from C18. The SPL used in this case is the relation $f_{\rm bar,500c}\propto\left(M_{\rm 500c}\right)^{\alpha_{\rm SPL}}$. Fig. 3 presents these results along with the baryon fraction of the universe measured by Planck Collaboration et al. (2016a). The top panels show that the observational data do not directly match TNG100 or TNG300, likely due to the AGN feedback model in IllustrisTNG, which efficiently enriches the intra-cluster medium gas of low-massive clusters and galaxy groups (Pop et al., 2022a). As a consequence, the baryon fraction increases at low mass values. However, this does not impact our analysis, as we focused on the behaviour of the scaling relation. We found that in IllustrisTNG, the relations between the baryon fraction and halo mass yielded a slope $\alpha_{\rm SPL}\simeq 0.248\pm 0.018$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.261\pm 0.006$ for TNG300. Without considering redshift, C18 data provides an SPL scaling relation with a slope of $\alpha_{\rm Chiu}\simeq 0.350\pm 0.076$. The scaling relations for TNG100, TNG300, and C18 data show similar behaviour, with baryon fractions increasing at high halo masses and decreasing at low halo masses. Furthermore, the slope values of IllustrisTNG and C18 agree within $1-2\sigma$. This agreement suggests that the IllustrisTNG results may provide insights into the missing baryons problem and help reconcile the tension with the hierarchical scenario highlighted by C18.
@@
In the bottom panels of Fig. (3), we show how the behaviour of the scaling relations changes when the remaining gas ($T<10^{7}$ K) and all remaining stellar mass in the halo are included. A direct comparison with C18 is not possible, as the parameters are not directly comparable. Although the general behaviour remains, showing an increase in the baryon fraction, the slopes differ significantly between simulations and observations. We found a slope $\alpha_{\rm SPL}\simeq 0.071\pm 0.010$ for TNG100 and $\alpha_{\rm SPL}\simeq 0.120\pm 0.003$ for TNG300 in the scaling relations. These lower slopes result from including the remaining gas and accounting for the ICL in the measurements, which benefits low-mass haloes more than high-mass ones. Consequently, low-mass haloes exhibit a greater fraction of cold gas compared to high-mass haloes (see Sections 4.2 and A). Additionally, more haloes in both simulations reach the universal baryon fraction. Most haloes that do not reach this value have masses of $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}\lesssim 14$. This mass threshold aligns with values reported in numerous studies discussing the breakdown of self-similarity in scaling relations (Pop et al., 2022a) and the redistribution of baryons in low-mass haloes by non-gravitational processes, such as AGN feedback (Ayromlou et al., 2023).